OpenStudy (anonymous):

Find the domain of these functions:

OpenStudy (anonymous):

\[f(x)=\frac{ \sqrt{x^2-4} }{ x^2-9 }\] \[g(x)=\ln(3x-2)\]

OpenStudy (anonymous):

@agent0smith @electrokid @RadEn

OpenStudy (raden):

for f(x) = sqrt(x^2-4)/(x^2-9) so that the real values, x^2 - 4 >= 0 solve for x also, the denominator cant be equal 0 it means x^2 - 9 not equal 0 solve for x

OpenStudy (anonymous):

so \[x \neq 0,2\]

OpenStudy (raden):

u just trying solve for the second case, right ? let it is equal = 0 so, x^2 - 9 = 0 factor out! (x+3)(x-3) = 0 the zeroes does satisfied for x = -3 or x = 3 so, x≠-3, 3

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

how about g(x)

OpenStudy (abb0t):

The domain is all the values a variable can take on without causing an invalid function. So what values can you plug into "x" for g(x) that will make it invalid.

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so domain is \[x \neq 0\]

OpenStudy (agent0smith):

@lilfayfay domain of the first function is not x≠0,2 reread RadEn's post: so that the real values, x^2 - 4 >= 0 solve for x also, the denominator cant be equal 0 it means x^2 - 9 not equal 0 solve for x

OpenStudy (anonymous):

yeah i got the first one, i dont get the 2nd one

OpenStudy (agent0smith):

For g(x)=ln(3x−2) 3x-2>0 that will give the domain

OpenStudy (anonymous):

so the domain is x>2/3

OpenStudy (agent0smith):

What was your domain for \[\Large f(x)=\frac{ \sqrt{x^2-4} }{ x^2-9 }\]

OpenStudy (anonymous):

\[x \neq \pm 3\]

OpenStudy (agent0smith):

"so the domain is x>2/3" correct for g(x)

OpenStudy (anonymous):

can you help me with another one please? ^^

OpenStudy (abb0t):

We will not do your homework for you :3

OpenStudy (agent0smith):

x≠±3 is not correct for f(x) \[\Large f(x)=\frac{ \sqrt{x^2-4} }{ x^2-9 }\] reread this (both parts): x^2 - 4 >= 0 solve for x also, the denominator cant be equal 0 it means x^2 - 9 not equal 0 solve for x

OpenStudy (anonymous):

but this is openstudy :(

OpenStudy (anonymous):

and it would really help prepare me for the AP Calculus test next week

OpenStudy (agent0smith):

You need BOTH of those parts for the domain of f.